I have a 5th grader that is trying to figure out what to do on his homework. the question is... Follow the pattern
281.8 x 1 = 281.8
281.8 x 10 = -------
281.8 x 100 = ------

Answer by CubeyThePenguin(3113)

281.8 * 1 = 281.8
281.8 * 10 = 2818.
281.8 * 100 = 28180.